OpenStudy (anonymous):

In there a difference between log(x-3) or log x-3 ?

OpenStudy (lgbasallote):

yes.. in log (x-3), you're "logging" (x-3). However, in log x - 3, you're just "logging" x. does that help?

OpenStudy (ash2326):

Yeah, \[(\log (x-3))\] \[(\log x)- 3\]

OpenStudy (anonymous):

yes it does thank you:)

OpenStudy (anonymous):

also..... compared to \(\large y=logx \) \(\large y=log(x-3) \) is a horizontal shift to the right 3 units. \(\large y=logx-3 \) is a vertical shift 3 units down.